If a person travels 40 mph going to city A to B, and 60 mph returning from B to A, what is the average speed of the entire trip? Explain.

I told my teacher that the average speed was 50 mph and got marked wrong. Where did I go wrong?

Answer by mananth(16949)

lets be practical
while going to B
speed = 40 mph
Assume it takes 1 hour
so distance = 40 miles
..
while returning the speed = 60 mph
distance = 40
time= 40/60 = 2/3 hours
..
Average speed for the trip = total distance / total time
..
Total distance = 80 miles
total time= 1+2/3 hours
=1 2/3
=5/3 hours
..
80/(5/3)
80*3/5
=16*3
=48 mph is the average speed
..
does your teacher approve of this logic
..
m.ananth@hotmail.ca